The gas-phase reaction of NONO with F2F2 to form NOFNOF and FF has an activation energy of EaEaEa = 6.30 kJ/molkJ/mol and a frequency factor of AAA1 = 6.00×108 M−1⋅s−1M−1⋅s−1 . The reaction is believed to be bimolecular:

NO(g)+F2(g)→NOF(g)+F(g)NO(g)+F2(g)→NOF(g)+F(g)

where NONO is a free radical. Free radicals, species with an odd number of unbound electrons, tend to be more reactive. This is one reason by NONO reacts so readily with gases such as fluorine or oxygen gas.

Part A:

What is the rate constant at 589 ∘C∘C ?

Express your answer to three significant digits with the appropriate units. For compound units, place a multiplication dot between units (e.g. J⋅mol−1⋅K−1J⋅mol−1⋅K−1).
Answer #1

Given:

T = 589.0 oC

=(589.0+273)K

= 862.0 K

A = 6*10^8 M-1.s-1

Ea = 6.3 KJ/mol

= 6300.0 J/mol

R = 8.314 J/mol.K

use:

K = A*e^(-Ea/RT)

= 6*10^8*e^(-6300.0/(8.314*862.0))

= 6*10^8*e^(-0.8791)

= 6*10^8*0.4152

= 2.491*10^8 M-1.s-1

Answer: 2.49*10^8 M-1.s-1